If "a" is a negative number find
/a/+a+/-2a/+/-a/
: If "a" is a negative number find
/a/+a+/-2a/+/-a/

Answer by josmiceli(2022)

Once you put a number inside the absolute value sign, it is made
positive no matter what it's sign is
abs(a) + a + abs(-2a) + abs(-a)
abs(a) + a + abs(-2a) + abs(-a) = a + a + 2a + a
a + a + 2a + a = 5a answer